# Chika Maduabuchi, Ericmoore Jossou, Matteo Bucci

### v1

@@ Abstract
Keywords HSV segmentation  $\cdot$ VideoSAM  $\cdot$ Phase detection
1 Introduction
@@ 2 Related Works
2.1 Traditional HSV Segmentation Methods
@@
2.2 Large Vision Foundation Models in Segmentation
@@
2.3 Existing Datasets in Boiling Phenomena
@@
2.4 Application of Large Vision Models to Scientific HSV Segmentation
@@
3 Methodology
3.1 Model Architecture
@@
Table 1: High-Speed Video Modalities for Experimental Data Collection
Modality Boiling Conditions Pressure (bar) Heat Flux (kW/m2) Wall Superheat (K) Size (px $\times$ px) Frames
Argon Saturated pool boiling 1 120 9.5 742$\times$564 6K
Nitrogen Saturated pool boiling 1 120 8.1 720$\times$602 6K
FC-72 Saturated pool boiling 1 170 32.6 386$\times$410 6K
Water Flow boiling at 500 kg/m2s 10 3000 11 676$\times$624 7.5K
3.2 Data Collection
To develop a versatile and generalizable foundation model for bubble segmentation in HSVs, we curated an extensive and diverse dataset of video frame-mask pairs from high-speed camera experiments. The dataset encompasses several boiling modalities with varied conditions, fluid properties, and imaging setups, ensuring robustness and applicability of the model across various scenarios. The HSV modalities are detailed in Table 1, which includes data collected under different boiling conditions for liquid argon, liquid nitrogen, FC-72, and high-pressure water. Each modality was recorded with a specific resolution and frame count to capture the dynamic boiling processes effectively.
3.3 Data Processing
@@ 3.6 Experimental Setup
3.6.1 Experiment 1: Zero-Shot Generalization Across Modalities
@@
3.6.2 Experiment 2: Performance Across Multiple Modalities
@@
3.6.3 Experiment 3: Comparison with U-Net CNN
@@
4 Results and Discussions
4.1 Experiment 1: Zero-Shot Generalization Across Modalities
@@
4.2 Experiment 2: Performance Across Multiple Modalities
@@
Fluid Model IoU F1 Score Precision
Water SAM 0.0620 0.1165 0.06
VideoSAM 0.1894 0.3143 0.18
FC-72 SAM 0.5721 0.7278 0.56
VideoSAM 0.7997 0.8885 0.83
Nitrogen SAM 0.6702 0.8025 0.65
VideoSAM 0.8317 0.9080 0.85
Argon SAM 0.6464 0.7852 0.64
VideoSAM 0.8384 0.9120 0.85
(a) Comparison of IoU, F1 Score, and Precision for VideoSAM and SAM.
(b) Box plots comparing the IoU and F1 Score of VideoSAM and SAM across different datasets.
Figure 4: Combined table and figure layout comparing the performance of VideoSAM across different datasets.
4.3 Experiment 3: Comparison with U-Net CNN
@@
Table 2 summarizes the IoU and F1 Scores for U-Net, VideoSAM, and SAM models across all datasets. The results clearly demonstrate VideoSAM’s superior performance in more complex environments, such as FC-72, Nitrogen, and Argon. For example, in the Argon dataset, VideoSAM achieved a mean IoU of 0.8384 and an F1 Score of 0.9120, outperforming both U-Net and SAM. Similarly, VideoSAM’s performance on the Nitrogen dataset (IoU: 0.8317, F1: 0.9080) also exceeded that of U-Net and SAM.
@@
Table 2: Comparison of IoU and F1 Score for U-Net, VideoSAM, and SAM Models across Different Fluids
Model Water FC-72 Nitrogen Argon
IoU F1 Score IoU F1 Score IoU F1 Score IoU F1 Score
U-Net 0.5619 0.7191 0.7244 0.8400 0.7547 0.8602 0.7815 0.8773
VideoSAM 0.1894 0.3143 0.7997 0.8885 0.8317 0.9080 0.8384 0.9120
SAM 0.0620 0.1165 0.5721 0.7278 0.6702 0.8025 0.6464 0.7852
The results from Table 2 indicate that while VideoSAM excels in handling more complex fluids, U-Net remains a strong contender in simpler datasets like Water. This suggests that the architecture of U-Net, particularly its cellular-level segmentation capabilities, makes it highly effective in environments with fewer objects and more consistent fluid structures, as seen in the Water dataset.
@@
4.4 Weaknesses and Potential Solutions
@@
5 Conclusion and Future Work

### v3

# VideoSAM: A Large Vision Foundation Model for High-Speed Video Segmentation

@@ Abstract
Index Terms:
Computer vision, Video segmentation, Segment Anything Model, Phase detection
I Introduction
@@ II Related Works
II-A Traditional HSV Segmentation Methods
@@
II-B Large Vision Foundation Models in Segmentation
@@
II-C Existing Datasets in Boiling Phenomena
@@
II-D Application of Large Vision Models to Scientific HSV Segmentation
@@
III Methodology
III-A Model Architecture
@@
Table I: High-Speed Video Modalities for Experimental Data Collection
Modality Conditions Heat Flux Frames
(kW/m2)
Argon SPB 120 6K
Nitrogen SPB 120 6K
FC-72 SPB 170 6K
Water FB (500 kg/m2s) 3000 7.5K
SPB: Saturated Pool Boiling, FB: Flow Boiling at 500 kg/m2s
III-B Data Collection
To develop a versatile and generalizable foundation model for bubble segmentation in HSVs, we curated an extensive and diverse dataset of video frame-mask pairs from high-speed camera experiments. The dataset encompasses several boiling modalities with varied conditions, fluid properties, pressures, and imaging setups, ensuring robustness and applicability across a wide range of scenarios. The HSV modalities are detailed in Table I, which includes data collected under saturated pool boiling (SPB) conditions at 1 bar for liquid argon, nitrogen, and FC-72. Additionally, high-pressure flow boiling (FB) experiments were conducted for water at 10 bar with a mass flux of 500 kg/m2s. Each modality was captured with specific resolutions and frame counts, effectively recording the dynamic behavior of these boiling processes.
III-C Data Processing
@@ III-F Experimental Setup
III-F1 Experiment 1: Zero-Shot Generalization Across Modalities
@@
III-F2 Experiment 2: Performance Across Multiple Modalities
@@
III-F3 Experiment 3: Comparison with U-Net CNN
@@
IV Results and Discussions
IV-A Experiment 1: Zero-Shot Generalization Across Modalities
@@
IV-B Experiment 2: Performance Across Multiple Modalities
@@
Fluid Model IoU F1 Score
Water U-Net 0.5619 0.7191
SAM 0.0620 0.1165
VideoSAM 0.1894 0.3143
FC-72 U-Net 0.7244 0.8400
SAM 0.5721 0.7278
VideoSAM 0.7997 0.8885
Nitrogen U-Net 0.7547 0.8602
SAM 0.6702 0.8025
VideoSAM 0.8317 0.9080
Argon U-Net 0.7815 0.8773
SAM 0.6464 0.7852
VideoSAM 0.8384 0.9120
(a) Comparison of IoU and F1 Score for U-Net, VideoSAM, and SAM across fluids.
(b) Box plots comparing IoU and F1 Score across datasets.
Figure 4: Combined table and figure layout comparing the performance of U-Net, VideoSAM, and SAM across different datasets.
IV-C Experiment 3: Comparison with U-Net CNN
@@
Table 4(a) summarizes the IoU and F1 Scores for U-Net, VideoSAM, and SAM models across all datasets. The results clearly demonstrate VideoSAM’s superior performance in more complex environments, such as FC-72, Nitrogen, and Argon. For example, in the Argon dataset, VideoSAM achieved a mean IoU of 0.8384 and an F1 Score of 0.9120, outperforming both U-Net and SAM. Similarly, VideoSAM’s performance on the Nitrogen dataset (IoU: 0.8317, F1: 0.9080) also exceeded that of U-Net and SAM.
@@
The results from Table 4(a) indicate that while VideoSAM excels in handling more complex fluids, U-Net remains a strong contender in simpler datasets like Water. This suggests that the architecture of U-Net, particularly its cellular-level segmentation capabilities, makes it highly effective in environments with fewer objects and more consistent fluid structures, as seen in the Water dataset.
@@
IV-D Weaknesses and Potential Solutions
@@
V Conclusion and Future Work